If 11.7 g of NH3 is isolated in a flask at a pressure of 1.75 bar, what will the density of the gas be at a temperature of 25C?
Answer #1

given

pressure (P) = 1.5 bar = 1.48 atm

molar mass of NH3(M) = 17 g/mole

T = 25 + 273 = 298 K

density (d) = ?

we know

density (d) = PM / RT = (1.48 * 17) / (0.0821 * 298) =1.03 g/lit